How much electricity (in kilowatt hours?) Would be produced by a generator / turbine if a 1 tonne weight was suspended from a cable and gravity was pulling it down for 100 meters? Or what would the formula be to solve this?
1
Expert's answer
2018-04-18T08:11:07-0400
All potential energy will transform into electricity.
E=mgh=1000*9.8*100=980000 J.
Now we need to express one kilowatt hour in joules: 1 kW*h=1000*3600 W*s=3600000 J.
Now we can express the amount of energy in kWh: E=980000/3600000=0.27 kWh.
Answer: 0.27 kWh.